# Text-Aware Diffusion for Policy Learning

### v1

@@ Abstract
Training an agent to achieve particular goals or perform desired behaviors is often accomplished through reinforcement learning, especially in the absence of expert demonstrations. However, supporting novel goals or behaviors through reinforcement learning requires the ad-hoc design of appropriate reward functions, which quickly becomes intractable. To address this challenge, we propose Text-Aware Diffusion for Policy Learning (TADPoLe), which uses a pretrained, frozen text-conditioned diffusion model to compute dense zero-shot reward signals for text-aligned policy learning. We hypothesize that large-scale pretrained generative models encode rich priors that can supervise a policy to behave not only in a text-aligned manner, but also in alignment with a notion of naturalness summarized from internet-scale training data. In our experiments, we demonstrate that TADPoLe is able to learn policies for novel goal-achievement and continuous locomotion behaviors specified by natural language, in both Humanoid and Dog environments. The behaviors are learned zero-shot without ground-truth rewards or expert demonstrations, and are qualitatively more natural according to human evaluation. We further show that TADPoLe performs competitively when applied to robotic manipulation tasks in the Meta-World environment.
@@ 1 Introduction
Our work looks towards natural language as a powerful interface through which humans can flexibly specify desired goals or behaviors of interest. We therefore investigate how to construct a zero-shot text-conditioned reward signal, replacing the need for ad-hoc designs, through which text-aligned policies can be learned. We present Text-Aware Diffusion for Policy Learning (TADPoLe), which utilizes a large-scale pretrained, frozen text-conditioned diffusion model to generate a dense reward signal for policy learning. We hypothesize that generative diffusion models, which are pretrained on internet-scale datasets to produce text-aligned, natural-looking images (saharia2022photorealistic, ; ramesh2022hierarchical, ) and videos (bar2024lumiere, ; girdhar2023emu, ; ho2022imagen, ), can be utilized to automatically craft a multimodal reward signal that encourages an agent to behave both faithfully with respect to text conditioning and naturally with respect to human perception. Our method is novel in its reward computation, as well as its utilization of a domain-agnostic generative model, rather than one trained from environment-specific or task-specific video demonstrations, as used in prior work torabi2018behavioral ; peng2021amp ; du2023learning ; escontrela2023video ; ma2023liv ; lynch2023interactive .
@@ 2 Related Work
Similar to our work, UniPi (du2023learning, ) treats the sequential decision-making problem as a text-conditioned video generation problem. The authors propose training a video diffusion model to produce a future visual plan for the agent; the subsequent frames are then converted to actions by means of an inverse dynamics model. VLP (du2023video, ) utilizes text-to-video generative models for planning and goal generation for an agent. Both methods require the video generative models to be trained ad-hoc on the target environments, whereas we directly use frozen general-purpose generative models. Mahmoudieh et al. (mahmoudieh2022zero, ) propose a framework that uses CLIP to generate a reward signal from a text description of a goal state and raw pixel observations from the environment, which is then used to learn a task policy. In VLM-RM (rocamonde2023vision, ), the authors also explore utilizing CLIP as the reward model for training humanoids to accomplish complex goal-reaching tasks. In our work, we investigate locomotion tasks on top of goal-achievement, and explore how using diffusion models to produce a reward signal can outperform CLIP-based approaches. Although not conditioned on text, VIPER  (escontrela2023video, ) also aims to harness recent advancements in generative modeling by employing a video prediction model’s likelihoods as a reward signal. However, VIPER does not enable the learning of policies conditioned on text, and requires in-domain expert videos for ad-hoc training the video model. Finally, Diffusion Reward (huang2023diffusion, ) also extracts a reward from a diffusion model to train policies; however, it requires training an ad-hoc video model on expert trajectories, the collection of which cannot always be assumed to be trivial, and does not enable text-conditioned policy learning.
@@ 3 Method
Figure 3: An illustration of the TADPoLe pipeline, which computes text-conditioned rewards through a pretrained, frozen diffusion model for policy learning. At each timestep, the subsequent frame rendered through the environment is corrupted with a sampled Gaussian source noise vector $\mathbf{\epsilon}_{0}$. The pretrained text-conditioned diffusion model then predicts the source noise that was added. The reward is designed to be large when the selected action produces frames well-aligned with the text prompt.
@@ 3.1 Text-Aware Diffusion for Policy Learning
We first describe how TADPoLe produces text-conditioned rewards from image observations. At each timestep $t$, reward $r_{t}$ is computed as a score between rendered subsequent image $\mathbf{o}_{t+1}$ and the provided text caption describing the behavior of interest, denoted by $y$, using a frozen, pretrained text-to-image diffusion model. We begin by corrupting the rendered image $\mathbf{o}_{t+1}$ with a sampled Gaussian source noise vector $\boldsymbol{\epsilon}_{0}\sim\mathcal{N}(\boldsymbol{\epsilon};\boldsymbol{0},\textbf{I})$ to produce noisy observation $\tilde{\mathbf{o}}_{t+1}$, and use the diffusion model to make an unconditional prediction $\mathbf{\hat{\epsilon}}_{\boldsymbol{\phi}}(\tilde{\boldsymbol{o}}_{t+1};\texttt{t}_{\text{noise}})$ as well as a conditional prediction $\mathbf{\hat{\epsilon}}_{\boldsymbol{\phi}}(\tilde{\boldsymbol{o}}_{t+1};\texttt{t}_{\text{noise}},y)$. Here $\mathbf{\hat{\epsilon}}_{\boldsymbol{\phi}}$ is a neural network that predicts the source noise given $\tilde{\boldsymbol{o}}_{t+1}$, the level of noise corruption $\texttt{t}_{\text{noise}}$, and optionally the text prompt $y$. We then compute the mean squared error (MSE) between the two predictions as a reward signal $r_{t}^{\text{align}}$ to be maximized:
@@
$r_{t}^{\text{rec}}=\left\lVert\boldsymbol{\hat{\epsilon}}_{\boldsymbol{\phi}}(\tilde{\boldsymbol{o}}_{t+1};\texttt{t}_{\text{noise}})-\boldsymbol{\epsilon}_{0}\right\rVert_{2}^{2}-\left\lVert\boldsymbol{\hat{\epsilon}}_{\boldsymbol{\phi}}(\tilde{\boldsymbol{o}}_{t+1};\texttt{t}_{\text{noise}},y)-\boldsymbol{\epsilon}_{0}\right\rVert_{2}^{2},$
@@
TADPoLe is agnostic to the specific choices of policy network architecture and optimization objectives. A pseudocode of the method is provided in Algorithm 1. It is worth emphasizing that $\texttt{t}_{\text{noise}}$ and subscript-less $t$ refer to different notions of time; $t$ indexes the timestep of the agent in the environment, whereas $\texttt{t}_{\text{noise}}$ determines the level of noise to corrupt the raw observed image.
@@
1:txt_caption = sample(action_phrase)
@@
11:               $\tilde{\boldsymbol{o}}_{t+1}\leftarrow\texttt{noisify}(\boldsymbol{o}_{t+1},\texttt{t}_{\text{noise}})$
@@
13:               $\tau\leftarrow\tau\cup{(\boldsymbol{s}_{t},\boldsymbol{a}_{t},\boldsymbol{r}_{t},\boldsymbol{s}_{t+1})}$
@@ 4.1 Experimental Setup and Evaluation
Benchmarks: We present our main results using the Dog and Humanoid environments from the DeepMind Control Suite tunyasuvunakool2020dm_control , and robotic manipulation tasks from Meta-World yu2020meta . Dog and Humanoid are known to be challenging due to their large action space, complex transition dynamics, and lack of task-specific priors (such as termination conditions). We update the environments by modifying the terrain rendered by MuJoCo todorov2012mujoco to have green grass and blue sky. We also limit the number of environment timesteps to be 300, which is sufficient to demonstrate successful learning of a behavior, rather than the default 1000. The agent’s initialized joint configurations were also fixed, as we focus on learning text-conditioned capabilities rather than robustness to initialization conditions. Meta-World is initially designed for multi-task and meta reinforcement learning, and is later adopted to evaluate language-conditioned imitation learning algorithms nair2022r3m ; mu2024embodiedgpt ; sontakke2024roboclip . We select a suite of tasks which are balanced for diversity and complexity, and each task is paired with a text prompt (see tasks and their corresponding prompts in Appendix C). Following prior design Huang2023DiffusionReward for Meta-World, we also add a sparse success signal to the dense text-conditioned reward signals.
@@
A main benefit of utilizing a reward signal conditioned flexibly on text is the ability to learn policies with behaviors beyond those defined by existing ground-truth reward functions. As these have no corresponding ground-truth reward functions, quantitative comparison across different text-conditioned methods is challenging; we therefore appeal to a qualitative user study. We perform a paid study through the Prolific platform, with a total of 25 anonymous random participants. For a video demonstration from each trained model, each participant is asked if it sufficiently aligns with the text prompt it was conditioned on. These results are depicted in tables as checkmarks () and x-marks (), where a checkmark denotes if a majority of participants believe it is text-aligned. We then proceed with a user study regarding naturalness. Given a video produced by VLM-RM and TADPoLe, users are given a choice as to which they believed to be the more natural motion or pose. This seeks to approximate how naturally the resulting Humanoid and Dog policies behave, according to human belief over how people and dogs naturally move in the real world.
@@
4.2 Goal Achievement
For text-conditioned goal-achievement, the objective is to learn a policy to consistently achieve a particular pose described by a text prompt; as the emphasis is for every frame to match a fixed goal pose rather than performing continuous motion, it is natural to apply TADPoLe with text-to-image diffusion models. We set the noise level $\texttt{t}_{\text{noise}}\sim{U}(400,500)$.
@@ 4.3 Continuous Locomotion
Table 3: Results for continuous locomotion experiments. For rows with an associated ground-truth reward function, numerical results are listed; for performant approaches, we report mean and standard deviation across 5 seeds. Evaluation for text-alignment is also reported. Video-TADPoLe greatly outperforms ViCLIP-RM and VLM-RM on both Humanoid and Dog.
@@
For the Humanoid task, we find that Video-TADPoLe achieves the best results amongst methods trained purely from visual and/or language feedback as in Table 3. On the other hand, ViCLIP-RM indeed learns to take steps, but does so sideways while maintaining an unnaturally lop-sided pose. VLM-RM also keeps the agent upright, but refrains from learning substantial movement, whereas LIV and Text2Reward fail to learn meaningful behavior.
For the Dog task, we notice that both VLM-RM as well as ViCLIP-RM collapse; they both learn to strike a particular pose and maintain it for perpetuity. Text2Reward, which does not have access to any visual information, but does have access to ground-truth state information for the Dog including speed, direction, and joint positions, achieves a reward of 63.15. Ultimately, Video-TADPoLe achieves a comparable result using a context window of 4, while also distinguishing itself as the most natural-looking policy in qualitative terms as the Dog agent appears to perform step-taking motions, rather than remain stationary. Table 2 further showcases a higher preference for the naturalness of the learned policies for continuous locomotion achieved by Video-TADPoLe compared to VLM-RM, in both Dog and Humanoid environments.
@@ 4.4 Robotic Manipulation
Success Rate (%) Door Open Door Close Drawer Open Drawer Close Window Open Window Close Coffee Push Button Press Soccer Lever Pull Average
VLM-RM 0 ($\pm$ 0) 79.7 ($\pm$ 39.8) 10.0 ($\pm$ 30.0) 100.0 ($\pm$ 0) 9.7 (±$\pm 29.0$) 0 ($\pm$ 0) 4.0 ($\pm$ 12.0) 30.0 ($\pm$ 45.8) 5.3 ($\pm$ 11.6) 11.3 ($\pm$ 18.3) 25.0
TADPoLe 40.0 ($\pm$ 49.0) 100.0 ($\pm$ 0) 45.3 ($\pm$ 46.6) 100.0 ($\pm$ 0) 74.0 ($\pm$ 37.9) 30.0 ($\pm$ 45.8) 18.6 ($\pm$ 20.5) 73.0 ($\pm$ 38.5) 25.0 ($\pm$ 15.4) 0 ($\pm$ 0) 50.6

### v2

@@ Abstract
Training an agent to achieve particular goals or perform desired behaviors is often accomplished through reinforcement learning, especially in the absence of expert demonstrations. However, supporting novel goals or behaviors through reinforcement learning requires the ad-hoc design of appropriate reward functions, which quickly becomes intractable. To address this challenge, we propose Text-Aware Diffusion for Policy Learning (TADPoLe), which uses a pretrained, frozen text-conditioned diffusion model to compute dense zero-shot reward signals for text-aligned policy learning. We hypothesize that large-scale pretrained generative models encode rich priors that can supervise a policy to behave not only in a text-aligned manner, but also in alignment with a notion of naturalness summarized from internet-scale training data. In our experiments, we demonstrate that TADPoLe is able to learn policies for novel goal-achievement and continuous locomotion behaviors specified by natural language, in both Humanoid and Dog environments. The behaviors are learned zero-shot without ground-truth rewards or expert demonstrations, and are qualitatively more natural according to human evaluation. We further show that TADPoLe performs competitively when applied to robotic manipulation tasks in the Meta-World environment, without having access to any in-domain demonstrations.
@@ 1 Introduction
Our work looks towards natural language as a powerful interface through which humans can flexibly specify desired goals or behaviors of interest. We therefore investigate how to construct a zero-shot text-conditioned reward signal, replacing the need for ad-hoc designs, through which text-aligned policies can be learned. We present Text-Aware Diffusion for Policy Learning (TADPoLe), which utilizes a large-scale pretrained, frozen text-conditioned diffusion model to generate a dense reward signal for policy learning. We hypothesize that generative diffusion models, which are pretrained on internet-scale datasets to produce text-aligned, natural-looking images (saharia2022photorealistic, ; ramesh2022hierarchical, ) and videos (bar2024lumiere, ; girdhar2023emu, ; ho2022imagen, ), can be utilized to automatically craft a multimodal reward signal that encourages an agent to behave both faithfully with respect to text conditioning and naturally with respect to human perception. Our method is novel in its reward computation, as well as its utilization of a domain-agnostic generative model, rather than one trained from environment-specific or task-specific video demonstrations, as used in prior work torabi2018behavioral ; peng2021amp ; du2024learning ; escontrela2024video ; ma2023liv ; lynch2023interactive .
@@ 2 Related Work
Similar to our work, UniPi (du2024learning, ) treats the sequential decision-making problem as a text-conditioned video generation problem. The authors propose training a video diffusion model to produce a future visual plan for the agent; the subsequent frames are then converted to actions by means of an inverse dynamics model. VLP (du2023video, ) utilizes text-to-video generative models for planning and goal generation for an agent. Both methods require the video generative models to be trained ad-hoc on the target environments, whereas we directly use frozen general-purpose generative models. Mahmoudieh et al. (mahmoudieh2022zero, ) propose a framework that uses CLIP to generate a reward signal from a text description of a goal state and raw pixel observations from the environment, which is then used to learn a task policy. In VLM-RM (rocamonde2023vision, ), the authors also explore utilizing CLIP as the reward model for training humanoids to accomplish complex goal-reaching tasks. In our work, we investigate locomotion tasks on top of goal-achievement, and explore how using diffusion models to produce a reward signal can outperform CLIP-based approaches. Although not conditioned on text, VIPER  (escontrela2024video, ) also aims to harness recent advancements in generative modeling by employing a video prediction model’s likelihoods as a reward signal. However, VIPER does not enable the learning of policies conditioned on text, and requires in-domain expert videos for ad-hoc training the video model. Finally, Diffusion Reward (huang2023diffusion, ) also extracts a reward from a diffusion model to train policies; however, it requires training an ad-hoc video model on expert trajectories, the collection of which cannot always be assumed to be trivial, and does not enable text-conditioned policy learning.
@@ 3 Method
Figure 3: An illustration of the TADPoLe pipeline, which computes text-conditioned rewards for policy learning through a pretrained, frozen diffusion model. At each timestep, the subsequent frame rendered through the environment is corrupted with a sampled Gaussian source noise vector $\mathbf{\epsilon}_{0}$. The pretrained text-conditioned diffusion model then predicts the source noise that was added. The reward is designed to be large when the selected action produces frames well-aligned with the text prompt.
@@ 3.1 Text-Aware Diffusion for Policy Learning
We first describe how TADPoLe produces text-conditioned rewards from image observations. At each timestep $t$, reward $r_{t}$ is computed as a score between rendered subsequent image $\mathbf{o}_{t+1}$ and the provided text caption describing the behavior of interest, denoted by $y$, using a frozen, pretrained text-to-image diffusion model. We begin by corrupting the rendered image $\mathbf{o}_{t+1}$ with a sampled Gaussian source noise vector $\boldsymbol{\epsilon}_{0}\sim\mathcal{N}(\boldsymbol{\epsilon};\boldsymbol{0},\textbf{I})$ to produce noisy observation $\tilde{\mathbf{o}}_{t+1}$, and use the diffusion model to make an unconditional prediction $\mathbf{\hat{\epsilon}}_{\boldsymbol{\phi}}(\tilde{\boldsymbol{o}}_{t+1};\texttt{t}_{\text{noise}})$ as well as a conditional prediction $\mathbf{\hat{\epsilon}}_{\boldsymbol{\phi}}(\tilde{\boldsymbol{o}}_{t+1};\texttt{t}_{\text{noise}},y)$. Here $\mathbf{\hat{\epsilon}}_{\boldsymbol{\phi}}(\cdot)$ is a neural network that predicts the source noise given $\tilde{\boldsymbol{o}}_{t+1}$, the level of noise corruption $\texttt{t}_{\text{noise}}$, and optionally the text prompt $y$; we overload notation to have $\mathbf{\hat{\epsilon}}_{\boldsymbol{\phi}}$ represent the source noise prediction in Figure 3. We then compute the mean squared error (MSE) between the two predictions as a reward signal $r_{t}^{\text{align}}$ to be maximized:
@@
$r_{t}^{\text{rec}}=\left\lVert\boldsymbol{\hat{\epsilon}}_{\boldsymbol{\phi}}(\tilde{\boldsymbol{o}}_{t+1};\texttt{t}_{\text{noise}})-\boldsymbol{\epsilon}_{0}\right\rVert_{2}^{2}-\left\lVert\boldsymbol{\hat{\epsilon}}_{\boldsymbol{\phi}}(\tilde{\boldsymbol{o}}_{t+1};\texttt{t}_{\text{noise}},y)-\boldsymbol{\epsilon}_{0}\right\rVert_{2}^{2}.$
@@
The choice of using symlog as a reward normalization technique is thoroughly studied in Section 4.5. TADPoLe is agnostic to the specific choices of policy network architecture and optimization objectives. A pseudocode of the method is provided in Algorithm 1. It is worth emphasizing that $\texttt{t}_{\text{noise}}$ and subscript-less $t$ refer to different notions of time; $t$ indexes the timestep of the agent in the environment, whereas $\texttt{t}_{\text{noise}}$ determines the level of noise to corrupt the raw observed image.
@@
1:prompt = sample(action_phrase)
@@
11:               $\tilde{\boldsymbol{o}}_{t+1}\leftarrow\texttt{noisify}(\boldsymbol{o}_{t+1},\boldsymbol{\epsilon}_{0},\texttt{t}_{\text{noise}})$
@@
13:               $\tau\leftarrow\tau\cup{(\boldsymbol{s}_{t},\boldsymbol{a}_{t},r_{t},\boldsymbol{s}_{t+1})}$
@@ 4.1 Experimental Setup and Evaluation
Benchmarks: We present our main results using the Dog and Humanoid environments from the DeepMind Control Suite tunyasuvunakool2020dm_control , and robotic manipulation tasks from Meta-World yu2020meta . Dog and Humanoid are known to be challenging due to their large action space, complex transition dynamics, and lack of task-specific priors (such as termination conditions). We update the environments by modifying the terrain rendered by MuJoCo todorov2012mujoco to have green grass and blue sky. We also limit the number of environment timesteps to be 300, which is sufficient to demonstrate successful learning of a behavior, rather than the default 1000. The agent’s initialized joint configurations are also fixed, as we focus on learning text-conditioned capabilities rather than robustness to initialization conditions. Meta-World was initially designed for multi-task and meta-reinforcement learning, and was later adopted to evaluate language-conditioned imitation learning algorithms nair2022r3m ; mu2024embodiedgpt ; sontakke2024roboclip . We select a suite of tasks which are balanced for diversity and complexity, and pair each task with a text prompt (see tasks and their corresponding prompts in Appendix C). Following prior design huang2023diffusion for Meta-World, we also add a sparse success signal to the dense text-conditioned reward signals.
@@
A main benefit of utilizing a reward signal conditioned flexibly on text is the ability to learn policies with behaviors beyond those defined by existing ground-truth reward functions. As these have no corresponding ground-truth reward functions, quantitative comparison across different text-conditioned methods is challenging; we therefore appeal to a qualitative user study. We perform a paid study through the Prolific platform, with a total of 25 anonymous random participants without prior training to estimate a general response from the human population. For a video demonstration from each trained model, selected as the last timestep of policy training without cherry-picking, each participant is asked if it sufficiently aligns with the text prompt it was conditioned on. These results are depicted in tables as checkmarks () and x-marks (), where a checkmark denotes if a majority of participants believe it is text-aligned. In Table A7, we provide the fine-grained user study results on what percentage of the users believe the video achieved by the policy is appropriately text-aligned. We then proceed with a user study regarding naturalness. Given a video produced by VLM-RM and TADPoLe, users are given a choice as to which they believed to be the more natural motion or pose. This seeks to approximate how naturally the resulting Humanoid and Dog policies behave, according to human belief over how people and dogs naturally move in the real world.
4.2 Goal Achievement
For text-conditioned goal-achievement, the objective is to learn a policy to consistently achieve a particular pose described by a text prompt; as the emphasis is for every frame to match a fixed goal pose rather than performing continuous motion, it is natural to apply TADPoLe with text-to-image diffusion models. We set the noise level $\texttt{t}_{\text{noise}}\sim{U}(400,500)$, with intuition provided in Appendix A.
@@ 4.3 Continuous Locomotion
Table 3: Results for continuous locomotion experiments. For rows with an associated ground-truth reward function, numerical results are listed; for performant approaches, we report mean and standard deviation across 5 seeds. Evaluation for text-alignment is also reported. Video-TADPoLe greatly outperforms ViCLIP-RM on both Humanoid and Dog.
@@
For the Humanoid task, we find that Video-TADPoLe achieves the best results amongst methods trained purely from visual and/or language feedback as in Table 3. On the other hand, ViCLIP-RM indeed learns to take steps, but does so sideways while maintaining an unnaturally lopsided pose. Meanwhile, LIV and Text2Reward fail to learn meaningful behaviors.
For the Dog task, we notice that the policy learned via ViCLIP-RM collapses; it learns to strike a particular pose and maintain it for perpetuity. Text2Reward, which does not have access to any visual information, but does have access to ground-truth state information for the Dog including speed, direction, and joint positions, achieves a reward of 63.15. Ultimately, Video-TADPoLe achieves a comparable result using a context window of 4, while also distinguishing itself as the most natural-looking policy in qualitative terms as the Dog agent appears to perform step-taking motions, rather than remain stationary. Table 2 further showcases a higher preference for the naturalness of the learned policies for continuous locomotion achieved by Video-TADPoLe compared to ViCLIP-RM, in both Dog and Humanoid environments.
@@ 4.4 Robotic Manipulation
Success Rate (%) Door Open Door Close Drawer Open Drawer Close Window Open Window Close
VLM-RM 0 ($\pm$ 0) 79.7 ($\pm$ 39.8) 10.0 ($\pm$ 30.0) 100.0 ($\pm$ 0) 9.7 (±$\pm 29.0$) 0 ($\pm$ 0)
TADPoLe 40.0 ($\pm$ 49.0) 100.0 ($\pm$ 0) 45.3 ($\pm$ 46.6) 100.0 ($\pm$ 0) 74.0 ($\pm$ 37.9) 30.0 ($\pm$ 45.8)
Success Rate (%) Coffee Push Button Press Soccer Lever Pull Average
VLM-RM 4.0 ($\pm$ 12.0) 30.0 ($\pm$ 45.8) 5.3 ($\pm$ 11.6) 11.3 ($\pm$ 18.3) 25.0
TADPoLe 18.6 ($\pm$ 20.5) 73.0 ($\pm$ 38.5) 25.0 ($\pm$ 15.4) 0 ($\pm$ 0) 50.6
Table 5: Quantitative results for TADPoLe and Video-TADPoLe with and without the symlog normalization operation, averaged over 3 seeds. Hyperparameters such as weights and noise level were kept the same across all experiments. We discover that not only does using the symlog transformation enable the highest empirical rewards, it also facilitates the reuse of hyperparameters across tasks, environments, and other diffusion models.
Environment Method Prompt SymLog Direct Scaling SymExp Min-Max Standardization
Humanoid TADPoLe “a person standing" 267.23 239.74 241.49 256.81 236.59
Humanoid Video-TADPoLe “a person walking" 226.29 4.58 3.68 61.31 134.66
Dog Video-TADPoLe “a dog walking" 81.22 35.30 9.46 6.15 5.05
4.5 Normalization Study
Of interest is what reward normalization technique is most performant for adjusting the raw computed alignment and reconstruction terms into a final reward used for policy learning. We investigate a variety of normalization strategies in a quantitative manner in Table 5, reusing parameters $w_{1}=2000$ and $w_{2}=200$ across experiments, the selection of which is detailed in Section B.3 and Table A4.
Apart from the symlog transformation, we also compare against using no additional normalization (denoted as “Direct Scaling”), and using symexp. We also compare against empirical normalization techniques. This includes min-max rescaling, where an empirically estimated minimum value is subtracted from the achieved reward, which is then divided by an empirically calculated min-max range, and rescaled to $[-1,1]$. This also includes standardization, which subtracts an empirically estimated mean from the achieved reward and then divides it by an empirically estimated standard deviation. We apply these techniques across Humanoid and Dog environments, for both TADPoLe and Video-TADPoLe. We discover that the symlog operation is the reward normalization strategy that achieves the best empirical results across robotic configurations, visual environments, and desired tasks, while reusing the same hyperparameter settings. We hypothesize that it helps to normalize the raw computed reward signals across diffusion models and environments to be roughly on the same scale. Indeed, we showcase how removing the symlog transformation, as well as using other normalization techniques, reduces consistent policy learning performance across environments and diffusion models with the same fixed hyperparameters. We further note that it does not require empirically estimated values, unlike min-max and standardization.